The supplement of an angle is forty more than the triple of the angle, find this angle and the supplement

Answer by checkley79(3341)

LET X BE ONE ANGLE.
THEN THE SUPPLEMENTARY ANGLE IS:
3X+40
X+3X+40=180
4X=180-40
4X=140
X=140/40
X=35 DEGREES FOR THE SMALLER ANGLE.
3*35+40=105+40=145 DEGREES FOR THE LARGER ANGLE.
PROOF:
35+145=180
180=180